OpenStudy (anonymous):

Solve the proportion to determine the value of x. 2x5 = x+33

OpenStudy (ddos_dragon):

2X5= 10 So the 33 has to add to something that will equal 10. -23 + 33= 10 X= -23

OpenStudy (ddos_dragon):

I am fairly certain that that is correct.

OpenStudy (anonymous):

it cant out wrong it is actually written like 2x/5=x+3/33

OpenStudy (anonymous):

so is it the same answer ?

OpenStudy (anonymous):

@Ddos_Dragon

OpenStudy (ddos_dragon):

Are you sure that correct? 2x/5=x+3/33 seems unlikely to be the equation.